# Can LLM-based Financial Investing Strategies Outperform the Market in Long Run?

### v2

@@
Abstract
Large Language Models (LLMs) have recently been leveraged for asset pricing tasks and stock trading applications, enabling AI agents to generate investment decisions from unstructured financial data. However, most evaluations of LLM timing-based investing strategies are conducted on narrow timeframes and limited stock universes, overstating effectiveness due to survivorship and data-snooping biases. We critically assess their generalizability and robustness by proposing FINSABER [^1] Data and code available at https://github.com/waylonli/FINSABER/., a backtesting framework evaluating timing-based strategies across longer periods and a larger universe of symbols. Systematic backtests over two decades and 100+ symbols reveal that previously reported LLM advantages deteriorate significantly under broader cross-section and over a longer-term evaluation. Our market regime analysis further demonstrates that LLM strategies are overly conservative in bull markets, underperforming passive benchmarks, and overly aggressive in bear markets, incurring heavy losses. These findings highlight the need to develop LLM strategies that are able to prioritise trend detection and regime-aware risk controls over mere scaling of framework complexity.
Can LLM-based Financial Investing Strategies Outperform the Market in Long Run?
Weixian Waylon Li1        Hyeonjun Kim2        Mihai Cucuringu3,4        Tiejun Ma1 1School of Informatics, University of Edinburgh 2Global Finance Research Center, Sungkyunkwan University 3Department of Mathematics, University of California Los Angeles 4Department of Statistics, University of Oxford {waylon.li, tiejun.ma}@ed.ac.uk
1 Introduction
Large language models (LLMs) are increasingly used in financial decision-making, especially for generating investment actions such as Buy, Hold, or Sell (Ding et al., 2024a; Fatouros et al., 2025). These so-called LLM timing-based investing strategies leverage LLMs’ ability to interpret historical and real-time data to autonomously trade. From sentiment-driven trading Zhang et al. (2024a) to sophisticated multi-agent systems Yu et al. (2023); Zhang et al. (2024b), a growing body of work has explored the potential of LLMs as autonomous financial agents.
@@ 1 Introduction
Method Type (Period, Symbols) Code
MarketSenseAI Sentiment Driven (1y3m, 100)
TradingGPT Multi Agents (N/A, N/A)
FinMem Multi Agents (6m, 5)
FinAgent Multi Agents (6m, 6)
FinRobot Multi Agents (N/A, N/A)
TradExpert Multi Agents (1y, 30)
FinCon Multi Agents (8m, 8)
TradingAgents Multi Agents (3m, 3)
MarketSenseAI 2.0 Multi Agents (2y, 100)
Table 1: Summary of current LLM-based investing strategies. Evaluation period and symbol coverage are combined for compactness.
While recent efforts such as Wang et al. (2025) and Hu et al. (2025) have begun to address backtesting of deep learning (DL)-based strategies and benchmarking of LLM-based time-series forecasting, a standard framework for rigorously evaluating LLM investing strategies remains absent. Moreover, these works do not effectively tackle the mitigation of biases in quantitative trading. To fill this gap, we introduce FINSABER, a comprehensive framework for benchmarking LLM timing-based investing strategies that supports longer backtesting periods, a broader and more diverse symbol universe, and explicit bias mitigation. Specifically, our main contributions are:
  1. 1.
@@ 2 Related Works
3 Definitions of Investing Strategies
@@
Timing-based strategies generate daily Buy ($+1$), Sell ($-1$), or Hold (00) signals based on market data such as prices, technical indicators, or model outputs. The objective is to capture short-term price movements through systematic trading rules.
@@
Selection-based strategies identify subsets of assets expected to outperform based on ranking signals. Assets are selected periodically using top-$k$ or thresholding. These strategies focus on cross-sectional alpha rather than timing individual trades.
4 Why Broader and Longer Evaluation Matters
@@
The Case for Broader and Longer Evaluation.
@@
5 FINSABER
As discussed in §4, existing evaluations of LLM-based investors suffer from survivorship bias, look-ahead bias, and data-snooping bias. These issues are largely due to limited evaluation periods and narrow stock selections. In this study, all subsequent findings and analyses are derived from our meticulously constructed backtesting framework, FINSABER [^2] Financial INvesting Strategy Assessment with Bias mitigation, Expanded time, and Range of symbols, which systematically addresses biases and is specifically designed to meet the practical needs of LLM-based strategies, including the integration of unstructured, multi-source data.
FINSABER comprises three core modules: (1) a multi-source data module, (2) a modular strategy base, and (3) a bias-aware two-step backtesting pipeline. Figure 1 illustrates the framework.
Figure 1: Overview of the FINSABER Backtest Framework. The central pipeline illustrates the backtesting process. The framework includes a Strategies Base Module (green), which covers both selection-based and timing-based strategies, and a Multi-source Data Module (yellow), integrating diverse financial data inputs.
@@ 6 Experiments
Specifically, our experiments include two parts: (1) Pitfalls of selective evaluation: Replicating previously reported results on select periods and symbols, then extending this evaluation period to demonstrate performance deterioration. (2) Fair and robust comparisons: Implementing systematic stock-selection methods to explicitly mitigate survivorship and data-snooping biases for fairer LLM assessments. We only consider go-long positions, aligning with current LLM strategies.
6.1 Pitfalls of Selective Evaluation
@@
We begin by replicating earlier evaluation setups that demonstrated the effectiveness of LLM investing strategies on TSLA, NFLX, AMZN, and MSFT during the previously reported period (6 October 2022 to 10 April 2023). Additionally, we incorporate broader benchmarks, including traditional rule-based, ML, and DL methods. Previous studies omit key details such as exact risk-free rates and transaction costs. Thus, we explicitly set a historical average risk-free rate of 0.03 and use Moomoo’s [^3] https://www.moomoo.com/ca/support/topic10_122?lang=en-us standard US commission fee (\$0.0049/share, minimum \$0.99/order), comparable to HSBC and TradeUp [^4] https://www.tradeup.com/pricing/detail platforms.
Symbol Type Strategy SPR $\uparrow$ CR $\uparrow$ MDD $\uparrow$ AV $\downarrow$
FinMem Selection (2022-10-06 to 2023-04-10)
TSLA Rule Based Buy and Hold -0.342 -20.483 -52.729 55.910
SMA Cross -0.293 -5.540 -18.517 38.602
WMA Cross 0.215 3.741 -18.492 42.062
ATR Band -0.595 -19.142 -39.599 42.161
Bollinger Bands -0.769 -24.747 -44.655 45.366
Turn of The Month 0.219 3.639 -11.642 31.042
Predictor ARIMA 0.601 15.007 -24.446 41.402
XGBoost 0.331 6.213 -35.374 37.729
RL A2C - - - -
DDPG - - - -
PPO -0.343 -20.425 -52.594 55.861
SAC - - - -
TD3 - - - -
LLM FinMem (GPT-4o-mini) 0.927 19.940 -30.144 48.638
FinMem (GPT-4o) 0.404 5.312 -36.351 54.434
FinMem (reported) 2.679 61.776 -10.800 46.865
FinAgent - - - -
NFLX Rule Based Buy and Hold 1.326 43.079 -20.184 41.523
SMA Cross -1.020 -8.285 -15.942 20.477
WMA Cross -0.803 -6.004 -14.290 19.826
ATR Band 0.150 2.992 -12.231 19.314
Bollinger Bands -0.558 -4.996 -13.244 16.754
Turn of The Month 0.559 8.383 -10.641 17.194
Predictor ARIMA 1.159 23.783 -15.043 25.749
XGBoost 0.770 10.134 -11.246 14.928
RL A2C - - - -
DDPG - - - -
PPO - - - -
SAC - - - -
TD3 1.325 42.872 -20.121 41.448
LLM FinMem (GPT-4o-mini) 1.704 32.549 -13.018 34.766
FinMem (GPT-4o) 0.896 16.244 -15.234 38.209
FinMem (reported) 2.017 36.449 -15.850 36.434
FinAgent 1.543 41.167 -20.417 51.030
AMZN Rule Based Buy and Hold -0.460 -13.250 -31.546 35.624
SMA Cross -0.420 -4.433 -18.910 27.084
WMA Cross -0.563 -6.121 -21.030 26.831
ATR Band 0.622 11.007 -15.842 23.272
Bollinger Bands -0.402 -7.105 -20.615 26.559
Turn of The Month -0.037 0.039 -14.892 20.722
Predictor ARIMA -0.225 -4.752 -20.046 26.899
XGBoost 1.955 42.468 -8.816 25.135
RL A2C -0.462 -13.186 -31.397 35.496
DDPG -0.462 -13.186 -31.397 35.496
PPO -0.576 -9.485 -22.761 24.169
SAC - - - -
TD3 -0.462 -13.186 -31.397 35.496
LLM FinMem (GPT-4o-mini) 0.297 2.800 -2.744 10.247
FinMem (GPT-4o) -0.968 -20.091 -31.164 40.896
FinMem (reported) 0.233 4.885 -22.929 42.658
FinAgent -1.108 -6.113 -9.317 13.257
MSFT Rule Based Buy and Hold 0.974 21.171 -14.192 28.327
SMA Cross 1.515 18.289 -8.746 20.821
WMA Cross 1.334 16.576 -8.883 21.503
ATR Band 1.036 12.979 -7.709 15.005
Bollinger Bands 2.115 31.619 -3.475 18.243
Turn of The Month -0.034 0.970 -11.955 15.097
Predictor ARIMA 2.245 44.777 -7.121 22.636
XGBoost 0.895 12.678 -10.734 16.721
RL A2C 0.973 21.059 -14.120 28.117
DDPG - - - -
PPO - - - -
SAC - - - -
TD3 - - - -
LLM FinMem (GPT-4o-mini) -0.554 -7.104 -14.588 25.969
FinMem (GPT-4o) 0.792 12.834 -13.555 33.884
FinMem (reported) 1.440 23.261 -14.989 32.562
FinAgent 1.252 21.438 -14.502 32.952
COIN Rule Based Buy and Hold 0.170 -12.729 -54.402 86.718
SMA Cross 0.566 15.215 -33.219 90.853
WMA Cross 0.517 12.655 -33.259 94.180
ATR Band 0.933 25.169 -22.906 38.716
Bollinger Bands -0.795 -24.371 -40.733 43.904
Turn of The Month 0.713 21.242 -21.522 49.357
LLM FinMem (GPT-4o-mini) -1.674 -8.196 -8.196 11.582
FinMem (GPT-4o) -0.601 -38.368 -60.210 94.212
FinMem (reported) 0.717 34.983 -35.753 89.752
FinAgent 0.313 -10.598 -56.641 111.434
Table 2: Backtest performance of benchmark strategies over the previously reported period (2022-10-06 to 2023-04-10) where LLM investing strategies were shown to be effective. “-” metrics indicate no trading activities were triggered. Top in red and second-best in blue.
Table 2 summarises these results. Our analysis reveals that LLM investors are not universally superior, even within their originally favoured setups. Specifically, FinMem shows significant outperformance only for TSLA, whereas traditional benchmarks remain competitive or superior for other symbols. These observations caution against overly optimistic interpretations from selective evaluations. FinAgent, another LLM-based method, performs comparably to FinMem on NFLX and MSFT but generally does not offer consistent improvements across the set. Additionally, we find that LLM-based strategies exhibit high annual volatility and substantial maximum drawdowns, indicating a high-risk profile. This observation underscores the importance of explicit risk assessments in evaluating such strategies.
Additionally, our results highlight that the choice of underlying language model dramatically impacts strategy performance. The original FinMem uses GPT-4; however, switching to GPT-4o or GPT-4o-mini significantly alters results. Interestingly, larger models (GPT-4o) do not consistently outperform smaller ones (GPT-4o-mini), suggesting that model size alone is not predictive of superior performance and that model selection itself may introduce data-snooping biases. Further evidence in Appendix D reinforces this instability, where even a slight two-month extension of the evaluation period results in substantial variation for LLM-based strategies.
@@
Symbol Type Strategy SPR $\uparrow$ STR $\uparrow$ AR $\uparrow$ MDD $\uparrow$ AV $\downarrow$
Selected 4 (2004-01-01 to 2024-01-01)
TSLA Rule Based Buy and Hold 0.631 0.915 37.767 -50.839 45.242
SMA Cross 0.011 0.106 16.352 -42.975 29.144
WMA Cross 0.218 0.375 19.743 -31.866 27.471
ATR Band 0.021 0.062 -0.277 -38.516 27.262
Bollinger Bands 0.193 0.294 4.279 -37.166 26.267
Trend Following 0.139 0.624 17.342 -13.508 14.750
Turn of The Month 0.207 0.352 7.868 -27.904 23.595
Predictor ARIMA 0.681 1.003 24.138 -30.450 27.612
XGBoost 0.142 0.370 10.877 -22.901 19.537
RL A2C 0.172 0.249 3.875 -27.367 22.890
DDPG 0.446 0.653 26.265 -21.626 21.892
PPO 0.469 0.663 28.189 -46.810 40.156
SAC 0.119 0.190 6.654 -11.042 9.902
TD3 0.417 0.604 23.336 -33.725 30.233
LLM FinMem 0.641 1.069 42.153 -34.234 35.030
FinAgent 0.206 0.649 38.591 -36.930 38.302
NFLX Rule Based Buy and Hold 0.622 0.952 23.919 -48.119 41.704
SMA Cross -0.126 0.015 1.219 -32.856 22.373
WMA Cross -0.091 -0.008 1.586 -34.378 23.275
ATR Band 0.118 0.273 1.018 -37.646 24.373
Bollinger Bands 0.075 0.381 0.284 -34.002 23.090
Trend Following -0.248 -0.094 3.433 -15.066 14.528
Turn of The Month 0.288 0.489 7.122 -21.646 17.168
Predictor ARIMA 0.659 1.035 19.022 -27.567 25.514
XGBoost 0.202 0.355 4.957 -21.301 17.302
RL A2C 0.171 0.243 4.359 -20.960 16.129
DDPG 0.211 0.340 7.370 -27.922 22.317
PPO 0.541 0.814 19.279 -39.615 33.630
SAC 0.186 0.285 8.397 -9.545 9.216
TD3 0.291 0.431 10.900 -21.451 19.304
LLM FinMem 0.293 0.622 12.566 -27.721 26.876
FinAgent -0.419 0.621 22.543 -20.466 22.838
AMZN Rule Based Buy and Hold 0.551 0.829 15.997 -36.842 30.859
SMA Cross -0.265 -0.247 0.582 -24.203 17.142
WMA Cross -0.144 -0.100 2.281 -22.867 16.329
ATR Band 0.453 1.014 5.639 -19.922 15.166
Bollinger Bands 0.020 0.126 0.900 -23.757 15.764
Trend Following -0.041 0.438 7.153 -10.265 12.342
Turn of The Month -0.033 -0.013 1.479 -20.472 15.726
Predictor ARIMA 0.339 0.504 7.523 -20.612 19.115
XGBoost -0.587 -0.366 1.200 -13.659 11.106
RL A2C 0.165 0.247 3.925 -14.841 11.654
DDPG 0.252 0.391 7.059 -10.153 9.306
PPO 0.505 0.767 13.831 -29.128 24.392
SAC 0.179 0.257 4.438 -14.093 11.665
TD3 0.382 0.597 11.738 -21.942 19.149
LLM FinMem 0.188 0.340 5.695 -28.296 24.786
FinAgent 0.364 0.663 12.699 -25.516 25.390
MSFT Rule Based Buy and Hold 0.461 0.620 11.237 -25.457 21.787
SMA Cross -0.172 -0.167 1.948 -16.610 12.051
WMA Cross -0.226 -0.233 0.827 -16.708 12.083
ATR Band 0.316 0.636 5.719 -11.893 10.881
Bollinger Bands -0.053 -0.028 1.578 -16.102 11.933
Trend Following -0.366 -0.205 1.966 -8.846 7.714
Turn of The Month -0.264 -0.343 -0.180 -14.309 10.438
Predictor ARIMA 0.304 0.466 8.207 -15.227 13.819
XGBoost 0.171 0.322 5.890 -10.523 10.335
RL A2C 0.279 0.380 7.478 -13.447 11.933
DDPG 0.300 0.403 7.838 -14.007 12.351
PPO 0.344 0.463 8.589 -16.697 14.410
SAC 0.216 0.288 5.329 -14.866 11.835
TD3 0.050 0.070 1.405 -9.491 6.648
LLM FinMem 0.203 0.293 4.567 -19.270 17.891
FinAgent 0.285 0.432 11.123 -18.596 18.863
Table 3: Backtest performance of benchmark strategies for previously reported LLM-selected symbols over an extended period (2004-01-01 or earliest available to 2024-01-01). Top in red and second-best in blue.
Table 3 summarises these extended period results. Crucially, extending the evaluation horizon significantly diminishes the perceived superiority of LLM investors. Over two decades, traditional strategies like Buy and Hold consistently rank among the top performers across most symbols. TSLA is the only case where LLM investors (FinMem, FinAgent) clearly lead in AR, while for NFLX, AMZN, and MSFT, Buy and Hold or other strategies match or outperform them. This suggests previously reported LLM advantages are likely short-lived, potentially hand-picked, and highly sensitive to the evaluation period.

### v3

@@
Abstract.
Large Language Models (LLMs) have recently been leveraged for asset pricing tasks and stock trading applications, enabling AI agents to generate investment decisions from unstructured financial data. However, most evaluations of LLM timing-based investing strategies are conducted on narrow timeframes and limited stock universes, overstating effectiveness due to survivorship and data-snooping biases. We critically assess their generalizability and robustness by proposing FINSABER [^1] Data and code available at https://github.com/waylonli/FINSABER., a backtesting framework evaluating timing-based strategies across longer periods and a larger universe of symbols. Systematic backtests over two decades and 100+ symbols reveal that previously reported LLM advantages deteriorate significantly under broader cross-section and over a longer-term evaluation. Our market regime analysis further demonstrates that LLM strategies are overly conservative in bull markets, underperforming passive benchmarks, and overly aggressive in bear markets, incurring heavy losses. These findings highlight the need to develop LLM strategies that are able to prioritise trend detection and regime-aware risk controls over mere scaling of framework complexity.
Automated trading, LLM investors, Backtest, Benchmark
copyright: noneconference: ; Li et al.; 2025journalyear: 2025ccs: General and reference Evaluationccs: General and reference Empirical studiesccs: Computing methodologies Intelligent agents
1. Introduction
Large language models (LLMs) are increasingly used in financial decision-making, especially for generating investment actions such as Buy, Hold, or Sell (Ding et al., 2024a; Fatouros et al., 2025). These so-called LLM timing-based investing strategies leverage LLMs’ ability to interpret historical and real-time data to autonomously trade. From sentiment-driven trading (Zhang et al., 2024a) to sophisticated multi-agent systems (Yu et al., 2023; Zhang et al., 2024b), a growing body of work has explored the potential of LLMs as autonomous financial agents.
@@ 1. Introduction
Method Eval Period Eval Symbols Code
MarketSenseAI 1 year 3 months 100
TradingGPT N/A N/A
FinMem 6 months 5
FinAgent 6 months 6
FinRobot N/A N/A
TradExpert 1 year 30
FinCon 8 month 8
TradingAgents 3 months 3
MarketSenseAI 2.0 2 years 100
Table 1. Summary of current LLM-based investing strategies.
While recent efforts such as Wang et al. (2025) and Hu et al. (2025) have addressed benchmarking for deep learning (DL)-based trading and LLM-based time-series forecasting, comprehensive evaluation of LLM-based investing strategies remains unaddressed. Separately, FinBen (Xie et al., 2024) provides a thorough FinLLM benchmark covering multiple tasks, including decision-making. However, as a broad FinLLM benchmark, FinBen’s backtesting still relies on a limited, hand-picked symbol set, which contains the aforementioned biases and lacks a professional backtesting pipeline or systematic comparison with traditional strategies. To fill this gap, we introduce FINSABER, a comprehensive framework for benchmarking LLM timing-based investing strategies that supports longer backtesting periods, a broader and more diverse symbol universe, and explicit bias mitigation. Specifically, our main contributions are:
  1. (1)
@@ 2. Related Works
3. Definitions of Investing Strategies
@@
Timing-based strategies generate daily Buy (+1+1), Sell (1-1), or Hold (0) signals based on market data such as prices and technical indicators. The objective is to capture short-term price movements through systematic trading rules.
@@
Selection-based strategies identify subsets of assets expected to outperform based on ranking signals. Assets are selected periodically using top-$k$ or thresholding. These strategies focus on cross-sectional alpha.
4. Biases and Robustness Challenges in Backtesting LLM Investors
@@
Bias-Mitigation Requires Broader and Longer Evaluation
@@
5. FINSABER
As discussed in §4, existing evaluations of LLM-based investors suffer from survivorship bias, look-ahead bias, and data-snooping bias. These issues are largely due to limited evaluation periods and narrow stock selections. In this study, all subsequent findings and analyses are derived from our meticulously constructed backtesting framework, FINSABER [^2] Financial INvesting Strategy Assessment with Bias mitigation, Expanded time, and Range of symbols, which systematically addresses biases and meets the practical needs of LLM-based strategies, including the integration of unstructured, multi-source data. FINSABER comprises three core modules: (1) a multi-source data module, (2) a modular strategies base, and (3) a bias-aware two-step backtesting pipeline. Figure 1 illustrates the framework.
Figure 1. Overview of the FINSABER Backtest Framework. The central pipeline illustrates the backtesting process. The framework includes a Strategies Base Module (green), which covers both selection-based and timing-based strategies, and a Multi-source Data Module (yellow), integrating diverse financial data inputs.
@@ 5. FINSABER
Strategies Base.
We incorporate a comprehensive collection of strategies across multiple paradigms to ensure robust benchmarking. The timing-based strategies include open-source LLM investors (FinMem (Yu et al., 2023), FinAgent (Zhang et al., 2024b)), traditional rule-based approaches (Buy and Hold, Moving Average Crossover, Bollinger Bands (Bollinger, 2002), Trend Following (Wilcox et al., 2005)), ML/DL forecaster-based methods (ARIMA, XGBoost), and RL-based strategies (A2C, PPO, TD3, SAC implemented via FinRL (Liu et al., 2022) framework). Selection-based strategies encompass random K selection, Momentum Factor Selection (based on past returns), Volatility Effect Selection (selecting low-volatility stocks), and the stocks selection agent from the FinCon (Yu et al., 2024) framework. This diverse strategy base enables comprehensive performance comparison across different methodological approaches while maintaining extensibility for custom implementations. More technical details of the strategies are available in Appendix B.
@@ 6. Experiments
Specifically, our experiments include two parts: (1) Pitfalls of selective evaluation: Replicating previously reported results on select periods and symbols, then extending this evaluation period to demonstrate performance deterioration. (2) Fair and robust comparisons: Implementing systematic stock-selection methods to explicitly mitigate survivorship and data-snooping biases for fairer LLM assessments. We only consider go-long positions, aligning with current LLM strategies. Technical details, including hyperparameter configurations, are provided in Appendix E.
6.1. Pitfalls of Selective Evaluation
Type Strategy TSLA NFLX AMZN MSFT
[…]
FinMem Selection (2022-10-06 to 2023-04-10)
Rule
Based
Buy and Hold -0.342 -20.483 -52.729 55.910 1.326 43.079 -20.184 41.523 -0.460 -13.250 -31.546 35.624 0.974 21.171 -14.192 28.327
SMA Cross -0.293 -5.540 -18.517 38.602 -1.020 -8.285 -15.942 20.477 -0.420 -4.433 -18.910 27.084 1.515 18.289 -8.746 20.821
WMA Cross 0.215 3.741 -18.492 42.062 -0.803 -6.004 -14.290 19.826 -0.563 -6.121 -21.030 26.831 1.334 16.576 -8.883 21.503
ATR Band -0.595 -19.142 -39.599 42.161 0.150 2.992 -12.231 19.314 0.622 11.007 -15.842 23.272 1.036 12.979 -7.709 15.005
Bollinger Bands -0.769 -24.747 -44.655 45.366 -0.558 -4.996 -13.244 16.754 -0.402 -7.105 -20.615 26.559 2.115 31.619 -3.475 18.243
Turn of The Month 0.219 3.639 -11.642 31.042 0.559 8.383 -10.641 17.194 -0.037 0.039 -14.892 20.722 -0.034 0.970 -11.955 15.097
Predictor ARIMA 0.601 15.007 -24.446 41.402 1.159 23.783 -15.043 25.749 -0.225 -4.752 -20.046 26.899 2.245 44.777 -7.121 22.636
XGBoost 0.331 6.213 -35.374 37.729 0.770 10.134 -11.246 14.928 1.955 42.468 -8.816 25.135 0.895 12.678 -10.734 16.721
RL A2C -0.201 -15.876 -52.642 56.172 1.262 36.760 -20.436 37.542 -0.093 -3.253 -24.042 30.903 1.166 24.804 -13.437 26.743
PPO -0.254 -18.223 -52.609 57.301 1.420 40.181 -18.036 35.170 -0.576 -9.485 -22.761 24.169 1.149 25.752 -14.444 28.503
SAC -0.320 -20.598 -53.614 57.665 1.325 42.872 -20.121 41.448 -0.440 -13.215 -32.145 36.533 1.004 22.304 -14.522 28.904
TD3 -0.343 -20.423 -52.592 55.859 1.325 42.872 -20.121 41.448 -0.440 -13.215 -32.145 36.533 0.973 21.026 -14.099 28.073
LLM FinMem (GPT-4o-mini) 0.927 19.940 -30.144 48.638 1.704 32.549 -13.018 34.766 0.297 2.800 -2.744 10.247 -0.554 -7.104 -14.588 25.969
FinMem (GPT-4o) 0.404 5.312 -36.351 54.434 0.896 16.244 -15.234 38.209 -0.968 -20.091 -31.164 40.896 0.792 12.834 -13.555 33.884
FinMem (reported) 2.679 61.776 -10.800 46.865 2.017 36.449 -15.850 36.434 0.233 4.885 -22.929 42.658 1.440 23.261 -14.989 32.562
FinAgent - - - - 1.543 41.167 -20.417 51.030 -1.108 -6.113 -9.317 13.257 1.252 21.438 -14.502 32.952
Table 2. Backtest performance over the previously reported period (2022-10-06 to 2023-04-10) where LLM investing strategies were shown to be effective. “-” metrics indicate no trading activities were triggered. Top in red and second-best in blue.
@@ 6.1. Pitfalls of Selective Evaluation
We begin by replicating earlier evaluation setups that demonstrated the effectiveness of LLM investing strategies on TSLA, NFLX, AMZN, and MSFT during the previously reported period (6 October 2022 to 10 April 2023). Additionally, we incorporate broader benchmarks, including traditional rule-based, ML, and DL methods. Previous studies omit key details such as exact risk-free rates and transaction costs. Thus, we set a historical average risk-free rate of 0.03 and use Moomoo’s [^3] https://www.moomoo.com/ca/support/topic10_122 standard US commission fee (\$0.0049/share, minimum \$0.99/order), comparable to HSBC and TradeUp [^4] https://www.tradeup.com/pricing/detail.
Table 2 summarises these results. Our analysis indicates that LLM investors are not universally superior, even in their preferred setups. Specifically, FinMem only consistently outperforms for TSLA, while traditional benchmarks remain competitive or superior for other symbols. These results caution against overly optimistic interpretations from selective evaluations. FinAgent, the other LLM-based method, performs similarly to FinMem on NFLX and MSFT but generally lacks consistent improvements across the set. Furthermore, LLM-based strategies exhibit high annual volatility and significant maximum drawdowns, indicating a high-risk profile. This highlights the necessity of explicit risk assessments when evaluating such strategies.
Further evidence in Appendix D supports the instability of short-period evaluations, where even a slight two-month extension of the evaluation period results in substantial variation for LLM-based strategies.
@@
Type Strategy TSLA NFLX
[…]
Rule
Based
Buy and Hold 0.630 0.915 37.767 -50.839 45.243 0.622 0.952 23.919 -48.119 41.703
SMA Cross 0.680 1.013 23.681 -23.707 24.680 0.087 0.160 5.514 -28.689 21.836
WMA Cross 0.664 0.955 21.158 -25.135 24.087 0.004 0.071 1.447 -32.409 23.074
ATR Band 0.022 0.066 -0.005 -38.536 26.609 0.186 0.377 2.202 -35.603 23.922
Bollinger Bands 0.193 0.294 4.282 -37.157 26.267 0.075 0.381 0.286 -34.002 23.088
Trend Following 0.815 1.356 36.289 -28.113 28.628 0.403 0.646 11.868 -29.179 25.368
Turn of The Month 0.207 0.353 7.872 -27.902 23.595 0.287 0.487 7.097 -21.646 17.166
Predictor ARIMA 0.681 1.003 24.138 -30.450 27.612 0.659 1.035 19.022 -27.567 25.514
XGBoost 0.142 0.370 10.877 -22.901 19.537 0.202 0.355 4.957 -21.301 17.302
RL A2C 0.172 0.249 3.875 -27.367 22.890 0.171 0.243 4.359 -20.960 16.129
PPO 0.469 0.663 28.189 -46.810 40.156 0.541 0.814 19.279 -39.615 33.630
SAC 0.119 0.190 6.654 -11.042 9.902 0.186 0.285 8.397 -9.545 9.216
TD3 0.417 0.604 23.336 -33.725 30.233 0.291 0.431 10.900 -21.451 19.304
LLM FinMem 0.641 1.069 42.153 -34.234 35.030 0.293 0.622 12.566 -27.721 26.876
FinAgent 0.206 0.649 38.591 -36.930 38.302 -0.419 0.621 22.543 -20.466 22.838
Type Strategy AMZN MSFT
[…]
Rule
Based
Buy and Hold 0.551 0.829 15.997 -36.842 30.860 0.461 0.620 11.238 -25.463 21.791
SMA Cross 0.057 0.205 3.896 -22.096 17.520 -0.263 -0.314 0.192 -17.656 11.840
WMA Cross 0.175 0.300 5.702 -19.309 17.178 -0.363 -0.437 -1.664 -19.075 11.932
ATR Band 0.443 0.998 5.452 -19.990 15.130 0.317 0.637 5.725 -11.893 10.885
Bollinger Bands 0.019 0.125 0.895 -23.757 15.763 -0.054 -0.029 1.578 -16.101 11.931
Trend Following 0.649 1.111 16.018 -19.120 20.130 0.205 0.321 5.438 -17.515 13.419
Turn of The Month -0.029 -0.009 1.534 -20.422 15.728 -0.263 -0.343 -0.177 -14.308 10.438
Predictor ARIMA 0.339 0.504 7.523 -20.612 19.115 0.304 0.466 8.207 -15.227 13.819
XGBoost -0.587 -0.366 1.200 -13.659 11.106 0.171 0.322 5.890 -10.523 10.335
RL A2C 0.165 0.247 3.925 -14.841 11.654 0.279 0.380 7.478 -13.447 11.933
PPO 0.505 0.767 13.831 -29.128 24.392 0.344 0.463 8.589 -16.697 14.410
SAC 0.179 0.257 4.438 -14.093 11.665 0.216 0.288 5.329 -14.866 11.835
TD3 0.382 0.597 11.738 -21.942 19.149 0.050 0.070 1.405 -9.491 6.648
LLM FinMem 0.188 0.340 5.695 -28.296 24.786 0.203 0.293 4.567 -19.270 17.891
FinAgent 0.364 0.663 12.699 -25.516 25.390 0.285 0.432 11.123 -18.596 18.863
Table 3. Backtest performance for previously reported LLM-selected symbols over an extended period (2004-01-01 or earliest available to 2024-01-01). Top in red and second-best in blue.
@@
Table 3 summarises these extended period results. Crucially, extending the evaluation horizon significantly diminishes the perceived superiority of LLM investors. Over two decades, traditional strategies like Buy and Hold consistently rank among the top performers across most symbols. TSLA is the only case where LLM investors (FinMem, FinAgent) clearly lead in AR, while for NFLX, AMZN, and MSFT, Buy and Hold or other strategies match or outperform them. This further supports that previously reported LLM advantages are likely short-lived, potentially hand-picked, and highly sensitive to the evaluation period.